Use the elimination method to solve the given system of equations. Please show all work
7x + 2y = 10
-7x + y = -16

Answer by jim_thompson5910(35256)

Add the equations

7x + 2y = 10
-7x + y = -16
--------------
0x + 3y = -6

So 3y = -6 ----> y = -2 (dividing both sides by 3)

-------------------------------------------------------


Now use y = -2 and any equation with x,y in it to find x


7x + 2y = 10


7x + 2(-2) = 10


7x - 4 = 10


7x = 10 + 4


7x = 14


x = 14/7


So the solutions are x = 2 and y = -2